# Minimizing Surrogate Losses for Decision-Focused Learning using Differentiable Optimization

### v1

@@ Abstract
Decision-focused learning (DFL) trains a machine learning (ML) model to predict parameters of an optimization problem, to directly minimize decision regret, i.e., maximize decision quality. Gradient-based DFL requires computing the derivative of the solution to the optimization problem with respect to the predicted parameters. However, for many optimization problems, such as linear programs (LPs), the gradient of the regret with respect to the predicted parameters is zero almost everywhere. Existing gradient-based DFL approaches for LPs try to circumvent this issue in one of two ways: (a) smoothing the LP into a differentiable optimization problem by adding a quadratic regularizer and then minimizing the regret directly or (b) minimizing surrogate losses that have informative (sub)gradients. In this paper, we show that the former approach still results in zero gradients, because even after smoothing the regret remains constant across large regions of the parameter space. To address this, we propose minimizing surrogate losses – even when a differentiable optimization layer is used and regret can be minimized directly. Our experiments demonstrate that minimizing surrogate losses allows differentiable optimization layers to achieve regret comparable to or better than surrogate-loss based DFL methods. Further, we demonstrate that this also holds for DYS-Net, a recently proposed differentiable optimization technique for LPs, that computes approximate solutions and gradients through operations that can be performed using feedforward neural network layers. Because DYS-Net executes the forward and the backward pass very efficiently, by minimizing surrogate losses using DYS-Net, we are able to attain regret on par with the state-of-the-art while reducing training time by a significant margin.
@@ 1 Introduction
Many real-world decision-making problems can be cast as combinatorial optimization problems. Some parameters of these optimization problems (e.g., production costs or travel times) are typically unknown due to uncertainty when the decisions are made. In the context of data-driven contextual optimization [27], the problems of predicting parameters of an optimization problem can be viewed as “predict-then-optimize” (PtO) problems, including two key steps—the prediction of the unknown parameters given the observed contextual information and the subsequent optimization using those predicted parameters. Decision-focused learning (DFL) [21] trains machine learning (ML) models to predict uncertain parameters by directly minimizing the decision loss, a measure of the quality of the decisions made using the predicted parameters.
@@
In this paper, we focus on the first category of DFL approaches. In this category, the (MI)LP is turned into a differentiable optimization problem as follows: the linear programming (LP) relaxation of the (MI)LP is first considered by removing the integrality constraints on the variables. Then, a quadratic regularizer is added to the objective to convert the LP into a quadratic program (QP). In practice, the smoothing strength is kept reasonably low to ensure that it does not overshadow the true objective of the original optimization problem. As mentioned earlier, approaches in this category minimize the decision loss on the training dataset by computing the Jacobian of the solution to the smoothed QP with respect to the predicted parameters using a differentiable solver (e.g., Cvxpylayers [1]). The motivation for minimizing decision loss on the training set follows the principle of empirical risk minimization (ERM) [32], which suggests that improvements on the training data should lead to lower decision loss on unseen test instances.
@@
Furthermore, training in DFL is computation-intensive. Both types of approaches require solving the (smoothed) (MI)LP for each training instance with the predicted parameters in each epoch to compute the decision loss. This poses significant scalability challenges, and increases training time. To ease the computational burden, McKenzie et al. [22] recently developed a fully-neural optimization layer, DYS-Net. Unlike Cvxpylayers, DYS-Net computes the solution to the quadratically regularized LP and the gradient of the solution using neural operations such as matrix vector multiplications. By doing so, DYS-Net offers the potential to significantly speed up DFL training. However, this potential has not been fully established because McKenzie et al. [22] use datasets where the true parameters are not observed, while most DFL benchmarks [30] assume otherwise. So, DYS-Net has not been compared to the state-of-the-art DFL methods.We empirically show that minimizing surrogate losses with DYS-Net yields test-time regret comparable to state-of-the-art DFL methods, which direct regret minimization cannot match due to the zero-gradient issue. In summary, this paper makes the following contributions:
@@ 2 Background
We consider PtO for MILPs , where the vector of cost parameters $\bm{y}$ is to be predicted using a vector of contextual information $\bm{\psi}$, correlated with $\bm{y}$. PtO problems comprise two steps: predicting unknown parameters given contextual information and solving the MILP with these predicted parameters. In PtO problems, an ML model, $\mathcal{M}_{\theta}$ (with trainable parameters $\theta$), is trained using past observation pairs $\{(\bm{\psi}_{i},\bm{y}_{i})\}_{i=1}^{N}$ to map $\bm{\psi}\rightarrow\bm{y}$. Due to their successful performance in many predictive tasks, neural networks are commonly used as the predictive model in PtO settings [34]. We denote the predicted cost produced by the neural network, as $\bm{\hat{y}}$, i.e., $\bm{\hat{y}}=\mathcal{M}_{\theta}(\bm{\psi})$.
A straightforward approach to the PtO problem is to train $\mathcal{M}_{\theta}$ minimize the prediction error between $\bm{y}$ and $\bm{\hat{y}}$. However, previous works [12, 19, 34] show that such a prediction-focused approach can produce suboptimal decision performance. By contrast, in DFL, the ML model is directly trained to optimize the decision loss, which reflects the quality of the resulting decisions. When only the parameters in the objective function are predicted, the decision loss of interest is typically the regret, which measures the suboptimality of a decision resulting from a prediction. In DFL, one can consider other decision losses (Appendix A), such as squared decision errors (SqDE) between $\bm{w}^{\star}$ and $\bm{w}$. Regret and SqDE can be written in the following form:
@@ 2.2.2 Contrastive Loss
Nevertheless, computation of $\mathcal{L}_{\mathit{SPO}^{+}}$ or $\mathcal{L}_{SCE}$ entails solving the MILP with the predicted $\bm{\hat{y}}$ for each training instance in every epoch leading towards a significant computational burden. Mandi et al. [19] propose to minimize the LP relaxation for computing $\mathcal{L}_{\mathit{SPO}^{+}}$. Mulamba et al. [23] address the computational issue by using solution caching instead of repeatedly solving the MILP. The CaVE technique, proposed by Tang and Khalil [31] for ILPs minimize a different surrogate loss function, which avoids solving the problem during training by minimizing the angle between the predicted cost vector and the ‘normal cone’ of the true optimal solution.
@@ 3.1 Surrogate Losses without Differentiable Optimization
s.t. $(\bm{\hat{y}}+\bm{\Delta})^{\top}\bm{w}^{\star}>(\bm{\hat{y}}+\bm{\Delta})^{\top}\bm{w}^{\prime}$
@@ 3.2 Surrogate Losses with Differentiable Optimization
The introduction of smoothing ensures that the solution transitions smoothly, rather than abruptly, near the original LP’s transition points. However, the solution of the smoothed QP remains unchanged, or changes very slowly, in regions that are not near the transition points, provided that the smoothing strength is kept low as illustrated in Figure 7(b) in Appendix B. So, in this region, $\frac{d\bm{w}^{\star}(\bm{\hat{y}})}{d\bm{\hat{y}}}$ is nearly zero. When regret is minimized, its derivative with respect to the $\bm{\hat{y}}$ takes the following form:
@@
To prevent the derivative from vanishing far from the transition points, in this paper, we argue in favour of minimizing a surrogate loss even when $\mathit{SqDE}$ or $\mathcal{{\mathit{Regret}}}$ can be directly minimized using a differentiable solver. For instance, when $\mathcal{L}_{\mathit{SPO}^{+}}$ is minimized, the derivative of it after smoothing with respect to the $\bm{\hat{y}}$ would be:
@@ 3.4 Impact of Differentiable Optimization on $\mathcal{L}_{SCE}$
Further, note that if we minimize $\mathcal{L}_{SCE}$ or $\mathcal{L}_{\mathit{SPO}^{+}}$ using a non-differentiable solver, $\frac{\partial\bm{w}^{\star}(\bm{y})}{\partial\bm{y}}$ cannot be computed and only the first part of the derivative would be used. So, in such cases, Eq. 14 and Eq. 15 would reduce to SPO+ and SCE subgradients, respectively. We also mentioned in Section 3.1 that when $\mathcal{L}_{SCE}^{\hat{y}}$ and $\mathcal{L}_{SCE}^{(\hat{y}-y)}$ are minimized with a non-differentiable solver, they yield the same gradient. Note that this is not the case when using a differentiable solver. In that case, the gradient of $\mathcal{L}_{SCE}^{\hat{y}}$ differs from $\mathcal{L}_{SCE}^{(\hat{y}-y)}$, as shown below:
@@ 5 Experimental Evaluation
We will consider the following three optimization problems in our experiments. We provide detailed descriptions and mathematical formulations in Appendix F. See Appendix H for hyperparameters and tabulated results.  [^1] Source code: https://github.com/JayMan91/DYS-NET-SCE.
@@ 5.2 Experimental Results
The previous experiments showed that minimizing surrogate losses leads to lower test regret compared to directly minimizing regret using differentiable optimization. We now compare the quality and runtime of minimizing surrogate losses with different kinds of solvers. We will minimize both $\mathcal{L}_{\mathit{SPO}^{+}}$ and $\mathcal{L}_{SCE}$ using two non-differentiable solvers – one solving the ILP formulation (for ILP problems) and one solving the LP relaxation; as well as two differentiable solvers: Cvxpylayers (an exact differentiable solver) and DYS-Net (an approximate, faster differentiable solver), which solve the ‘smoothed’ QP. In Figure 4, for SP, we do not consider the ILP formulation, as the LP formulation, itself, provides the exact solution.
@@
To summarize, $SCE^{\text{DYS}}$ consistently matches or outperforms state-of-the-art DFL techniques, such as $SPO_{+}$ and PFY, in terms of regret, and significantly reduces training time across all problem instances – often by a factor of three or more compared to these methods. These results demonstrate that $SCE^{\text{DYS}}$ achieves low regret with faster training times, making it an efficient choice for DFL involving (MI)LPs.

### v2

@@ Abstract
Decision-focused learning (DFL) trains a machine learning (ML) model to predict parameters of an optimization problem, to directly minimize decision regret, i.e., maximize decision quality. Gradient-based DFL requires computing the derivative of the solution to the optimization problem with respect to the predicted parameters. However, for many optimization problems, such as linear programs (LPs), the gradient of the regret with respect to the predicted parameters is zero almost everywhere. Existing gradient-based DFL approaches for LPs try to circumvent this issue in one of two ways: (a) smoothing the LP into a differentiable optimization problem by adding a quadratic regularizer and then minimizing the regret directly or (b) minimizing surrogate losses that have informative (sub)gradients. In this paper, we show that the former approach still results in zero gradients, because even after smoothing the regret remains constant across large regions of the parameter space. To address this, we propose minimizing surrogate losses, even when a differentiable optimization layer is used and regret can be minimized directly. Our experiments demonstrate that minimizing surrogate losses allows differentiable optimization layers to achieve regret comparable to or better than surrogate-loss based DFL methods. Further, we demonstrate that this also holds for DYS-Net, a recently proposed differentiable optimization technique for LPs, that computes approximate solutions and gradients through operations that can be performed using feedforward neural network layers. Because DYS-Net executes the forward and the backward pass very efficiently, by minimizing surrogate losses using DYS-Net, we are able to attain regret on par with the state-of-the-art while reducing training time by a significant margin.
@@ 1 Introduction
Many real-world decision-making problems can be cast as combinatorial optimization problems. Some parameters of these optimization problems (e.g., production costs or travel times) are typically unknown due to uncertainty when the decisions are made. In the context of data-driven contextual optimization [27], the problems of predicting parameters of an optimization problem can be viewed as “predict-then-optimize” (PtO) problems, including two key steps – the prediction of the unknown parameters given the observed contextual information and the subsequent optimization using those predicted parameters. Decision-focused learning (DFL) [21] trains machine learning (ML) models to predict uncertain parameters by directly minimizing the decision loss, a measure of the quality of the decisions made using the predicted parameters.
@@
In this paper, we focus on the first category of DFL approaches. In this category, the (MI)LP is turned into a differentiable optimization problem as follows: the linear programming (LP) relaxation of the (MI)LP is first considered by removing the integrality constraints on the variables. Then the LP is converted into ‘smooth’ quadratic program (QP) by adding a quadratic regularizer to the objective function. In practice, the weight on the regularizer is kept low so that the smoothing strength remains small and does not overshadow the original objective. As mentioned earlier, approaches in this category minimize the decision loss on the training dataset by smoothing the LP into a QP and then computing the Jacobian of its solution with respect to the predicted parameters using a differentiable solver (e.g., Cvxpylayers [1]). The motivation for minimizing decision loss on the training set follows the principle of empirical risk minimization (ERM) [32], which suggests that improvements on the training data should lead to lower decision loss on unseen test instances.
@@
Furthermore, training in DFL is computation-intensive. Both types of approaches require solving the (smoothed) (MI)LP for each training instance with the predicted parameters in each epoch to compute the decision loss. This poses significant scalability challenges, and increases training time. To ease the computational burden, McKenzie et al. [22] recently developed a fully-neural optimization layer, DYS-Net. Unlike Cvxpylayers, DYS-Net computes the solution to the quadratically regularized LP and the gradient of the solution using neural operations such as matrix vector multiplications. By doing so, DYS-Net offers the potential to significantly speed up DFL training. However, this potential has not been fully established because McKenzie et al. [22] use datasets where the true parameters are not observed, while most DFL benchmarks [30] assume otherwise. So, DYS-Net has not been compared to the state-of-the-art DFL methods in this setting.We empirically show that our proposal of minimizing surrogate losses with DYS-Net yields test-time regret comparable to state-of-the-art DFL methods, which direct regret minimization cannot match due to the zero-gradient issue. In summary, this paper makes the following contributions:
@@ 2 Background
We consider PtO for MILPs , where the vector of cost parameters $\bm{y}$ is to be predicted using a vector of contextual information, $\bm{\psi}$, correlated with $\bm{y}$. PtO problems comprise two steps: predicting unknown parameters given contextual information and solving the MILP with these predicted parameters. In PtO problems, an ML model, $\mathcal{M}_{\theta}$ (with trainable parameters $\theta$), is trained using past observation pairs $\{(\bm{\psi}_{i},\bm{y}_{i})\}_{i=1}^{N}$ to map $\bm{\psi}\rightarrow\bm{y}$. Due to their successful performance in many predictive tasks, neural networks are commonly used as the predictive model in PtO settings [34]. We denote the predicted cost produced by the neural network, as $\bm{\hat{y}}$, i.e., $\bm{\hat{y}}=\mathcal{M}_{\theta}(\bm{\psi})$.
A straightforward approach to the PtO problem is to train $\mathcal{M}_{\theta}$ to minimize the prediction error between $\bm{y}$ and $\bm{\hat{y}}$. However, previous works [12, 19, 34] show that such a prediction-focused approach can produce suboptimal decision performance. By contrast, in DFL, the ML model is directly trained to optimize the decision loss, which reflects the quality of the resulting decisions. When only the parameters in the objective function are predicted, regret, which measures the suboptimality of a decision resulting from a prediction, is the decision loss of interest. In DFL, one can consider other decision losses (Appendix A), such as squared decision errors (SqDE) between $\bm{w}^{\star}$ and $\bm{w}$. Regret and SqDE can be written in the following form:
@@ 2.2.2 Contrastive Loss
Nevertheless, computation of $\mathcal{L}_{\mathit{SPO}^{+}}$ or $\mathcal{L}_{SCE}$ entails solving the MILP with the predicted $\bm{\hat{y}}$ for every training instance in each epoch, which creates a significant computational burden. Mandi et al. [19] propose to minimize the LP relaxation for computing $\mathcal{L}_{\mathit{SPO}^{+}}$. Mulamba et al. [23] address the computational issue by using solution caching instead of repeatedly solving the MILP. The CaVE technique, proposed by Tang and Khalil [31] for ILPs, minimizes a different surrogate loss, which measures the angle between the predicted cost vector and the ‘normal cone’ of the true optimal solution.
@@ 3.1 Surrogate Losses without Differentiable Optimization
$\text{s.t.}\quad(\bm{\hat{y}}+\bm{\Delta})^{\top}\bm{w}^{\star}>(\bm{\hat{y}}+\bm{\Delta})^{\top}\bm{w}^{\prime}$
@@ 3.2 Surrogate Losses with Differentiable Optimization
The introduction of smoothing ensures that the solution transitions smoothly, rather than abruptly, near the original LP’s transition points. However, the solution of the smoothed QP remains unchanged, or changes very slowly, in regions that are not near the transition points, provided that the smoothing strength is kept low as illustrated in Figure 7(b) in Appendix B. So, in this region, $\frac{d\bm{w}^{\star}(\bm{\hat{y}})}{d\bm{\hat{y}}}$ is nearly zero. When regret is minimized, its derivative with respect to $\bm{\hat{y}}$ takes the following form:
@@
To prevent the derivative from vanishing far from the transition points, in this paper, we argue in favour of minimizing a surrogate loss even when $\mathit{SqDE}$ or $\mathcal{{\mathit{Regret}}}$ can be directly minimized using a differentiable solver. For instance, when $\mathcal{L}_{\mathit{SPO}^{+}}$ is minimized, the derivative of it after smoothing with respect to $\bm{\hat{y}}$ would be:
@@ 3.4 Impact of Differentiable Optimization on $\mathcal{L}_{SCE}$
Furthermore, in Section 3.1, we showed that $\mathcal{L}_{SCE}^{\hat{y}}$ and $\mathcal{L}_{SCE}^{(\hat{y}-y)}$ produces the same gradient with a non-differentiable solver. In contrast, with a differentiable solver, the gradient of $\mathcal{L}_{SCE}^{\hat{y}}$ differs from that of $\mathcal{L}_{SCE}^{(\hat{y}-y)}$, as shown below:
@@
Also, note that if $\mathcal{L}_{SCE}$ or $\mathcal{L}_{\mathit{SPO}^{+}}$ is minimized with a non-differentiable solver, $\frac{\partial\bm{w}^{\star}(\bm{y})}{\partial\bm{y}}$ cannot be computed, so Eq.14 and Eq.15 reduce to SPO+ and SCE subgradients, respectively.
@@ 5 Experimental Evaluation
We will consider the following three optimization problems in our experiments. We provide detailed descriptions and mathematical formulations in Appendix F. See Appendix H for hyperparameters and tabulated results. The source code of our implementation is publicly available at https://github.com/JayMan91/DYS-NET-SCE.
@@ 5.2 Experimental Results
Previous experiments showed that, when using differentiable optimization, minimizing surrogate losses yields lower test regret than directly minimizing regret. We now compare the quality and runtime of minimizing surrogate losses with different kinds of solvers. We will minimize both $\mathcal{L}_{\mathit{SPO}^{+}}$ and $\mathcal{L}_{SCE}$ using two non-differentiable solvers – one solving the ILP formulation (for ILP problems) and one solving the LP relaxation; as well as two differentiable solvers: Cvxpylayers (an exact differentiable solver) and DYS-Net (an approximate, faster differentiable solver), which solve the ‘smoothed’ QP. In Figure 4, for SP, we do not consider the ILP formulation, as the LP formulation, itself, provides the exact solution.
@@
To summarize, $SCE^{\text{DYS}}$ consistently matches or outperforms state-of-the-art DFL techniques, such as $SPO_{+}$ and PFY, in terms of regret, and significantly reduces training time across all problem instances, often by a factor of three or more compared to these methods. These results demonstrate that $SCE^{\text{DYS}}$ achieves low regret with faster training times, suggesting that it is an efficient method for DFL.